A playground, 40 meters wide by 60 meters long, is to be doubled in are by extending each side an equal amount. By how much should each side be extended.

Answer by colliefan(242)

The original area is:
A = 40 m * 60 m = 2400 m^2
The desired new area is to be twice this or 4800 m^2.
We want to know what two new lengths we could use that when multiplied together gives the result 4800. If the previous numbers are going to be increased by 20 m, we are looking for a relatively even number. Consider factors of 48:
1 * 48 = 48 (so 10 * 4800 = 4800)
2 * 24 = 48 (so 20 * 2400 = 4800)
4 * 12 = 48 (so 40 * 120 = 4800)
6 * 8 = 48 (so 60 * 80 = 4800)
Which of these lengths meets the criteria?
If the lengths are 60 m and 80 m, they are 20 m longer than the originals.